# Quantum nonlinear effects in the number-conserving analogue gravity of Bose-Einstein condensates

### v2

@@ III.1 Linearized fluctuations around a background
and so on. For later convenience, we can thus establish a one-to-one correspondence with the subscript of each component of the density perturbation with the relevant scaling with $N^{1/2}$. For example, the subscript $0\rightarrow\mathcal{O}(N^{1/2})$ and similarly for other terms in descending order, i.e., an index $\alpha$ implies being of ${\cal O}(N^{(1-\alpha)/2})$. Note also that $f_{1}=\mathcal{O}(N^{0})$ and $f_{3}=\mathcal{O}(N^{-1})$ defined earlier in (LABEL:f1f2def) follow the same scaling convention, but that the phase perturbation index in (25) does not indicate a scaling with powers of $N^{1/2}$.
@@
The components of the metric $g_{\mu\nu}$ depend on the background variables $\rho_{0},\theta_{0}$, where the speed of sound in this background is $c^{2}_{s0}=U\rho_{0}$.
@@ III.2 Second-order fluctuations around a background
$\partial_{t}(\rho_{0}+\epsilon\delta\rho_{1})+\nabla\cdot\Big{(}(\rho_{0}+\epsilon\delta\rho_{1})(\nabla\theta_{0}+\epsilon\nabla\delta\theta_{1})\Big{)}=f_{1}(\rho_{0},\theta_{0})+\epsilon\left.\frac{\partial f_{1}}{\partial\rho_{c}}\right|_{0}\delta\rho_{1}+\epsilon\left.\frac{\partial f_{1}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}+\epsilon^{2}\left.\frac{\partial^{2}f_{1}}{\partial\rho_{c}\partial\theta_{c}}\right|_{0}\delta\rho_{1}\delta\theta_{1},$
$\partial_{t}(\theta_{0}+\epsilon\delta\theta_{1})+U(\rho_{0}+\epsilon\delta\rho_{1})+\frac{1}{2}(\nabla(\theta_{0}+\epsilon\delta\theta_{1}))^{2}+V+f_{3}(\rho_{0},\theta_{0})+\epsilon\left.\frac{\partial f_{3}}{\partial\rho_{c}}\right|_{0}\delta\rho_{1}+\epsilon\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}+\epsilon^{2}\left.\frac{\partial^{2}f_{3}}{\partial\rho_{c}\partial\theta_{c}}\right|_{0}\delta\rho_{1}\delta\theta_{1}=0.$
where we used the abbreviation $\left.\cdots\right|_{0}\coloneqq\left.\cdots\right|_{\rho_{0},\theta_{0}}$ here and in what follows. As we will argue later in detail, the two above equations define the new background on which we will consider the propagation of nonlinear fluctuations. Then the new equations governing the perturbations are
$\epsilon^{2}\partial_{t}\delta\rho_{2}+\nabla\cdot\Big{(}\epsilon^{2}\delta\rho_{2}(\nabla\theta_{0}+\epsilon\nabla\delta\theta_{1})+\epsilon^{2}(\rho_{0}+\epsilon\delta\rho_{1})\nabla\delta\theta_{2}\Big{)}=\epsilon^{2}\left.\frac{\partial^{2}f_{1}}{\partial\rho_{c}^{2}}\right|_{0}\delta\rho_{2}+\epsilon^{2}\left.\frac{\partial^{2}f_{1}}{\partial\theta_{c}^{2}}\right|_{0}\delta\theta_{2},$ (30)
$\epsilon^{2}\partial_{t}\delta\theta_{2}+\epsilon^{2}U\delta\rho_{2}+\epsilon^{2}\nabla(\theta_{0}+\epsilon\delta\theta_{1})\cdot\nabla\delta\theta_{2}+\epsilon^{2}\left.\frac{\partial^{2}f_{3}}{\partial\rho_{c}^{2}}\right|_{0}\delta\rho_{2}+\epsilon^{2}\left.\frac{\partial^{2}f_{3}}{\partial\theta_{c}^{2}}\right|_{0}\delta\theta_{2}=0.$ (31)
@@
where we truncated in the second line the series in Eqs. (24) and (25) after second order in $\epsilon$. We assume that these combinations satisfy the dynamical fluid Eqs. (21) and (22)) for two different boundary conditions, such that the expansions above are convergent, which in turn requires $\epsilon^{2}\delta\rho_{2}\ll\epsilon\delta\rho_{1}\ll\rho_{0}$. This also implies that we are not separately demanding $\rho_{0},\theta_{0}$ to follow any dynamical equations, at least formally. Then the condition on $\rho_{0},\theta_{0}$ is simply that they are two sufficiently slowly varying “background” functions, which are not necessarily solutions to the fluid dynamical equations, but can be chosen, for example, according to the ease of interpretation of experiments. Namely, such that the first and second order perturbative corrections on top of the background satisfy the fluid-dynamical equations.
@@
With respect to the variables ($\rho_{b},\theta_{b}$) of Eq. (32) defining the new background, we can substitute $\delta\rho_{2}$ from (31) into (30) to obtain
$\partial_{t}\Big{(}\frac{1}{\tilde{U}}\Big{[}\partial_{t}\delta\theta_{2}+{\bm{v}}_{b}\cdot\nabla\delta\theta_{2}+f^{\prime\prime}_{3}\delta\theta_{2}\Big{]}\Big{)}+\nabla\cdot\Bigg{(}\frac{1}{\tilde{U}}{\bm{v}}_{b}\Big{[}\partial_{t}\delta\theta_{2}+{\bm{v}}_{b}\cdot\nabla\delta\theta_{2}+f^{\prime\prime}_{3}\delta\theta_{2}\Big{]}-\rho_{b}\nabla\delta\theta_{2}\Bigg{)}\\ -\frac{\ddot{f}_{1}}{\tilde{U}}\Big{(}\partial_{t}\delta\theta_{2}+{\bm{v}}_{b}\cdot\nabla\delta\theta_{2}+f^{\prime\prime}_{3}\delta\theta_{2}\Big{)}+f^{\prime\prime}_{1}\delta\theta_{2}=0.$ (34)
Here, we have indicated partial derivatives with respect to $\rho_{c}$, and $\theta_{c}$ as overdot and prime, respectively. Furthermore, we have defined the background velocity
@@
$\tilde{U}=U+\left.\frac{\partial^{2}f_{3}}{\partial\rho_{c}^{2}}\right|_{0}.$ (36)
Note that $U$ and $f_{3}$ both scale as ${\cal O}(N^{-1})$.
@@ III.3 Lorentz-invariant wave equation for second-order perturbations
$\frac{1}{\sqrt{-\mathbf{g}}}\mathcal{D}_{\mu}\left(\sqrt{-\mathbf{g}}\mathbf{g}^{\mu\nu}\mathcal{D}_{\nu}\delta\theta_{2}\right)\\ -\mathbf{g}^{\mu\nu}\mathcal{A}_{\mu}\mathcal{D}_{\nu}\delta\theta_{2}-\mathcal{B}^{\mu}\mathcal{D}_{\mu}\delta\theta_{2}+\mathcal{M}^{2}\delta\theta_{2}=0.$ (37)
@@
$\mathbf{g}_{tt}=-\sqrt{\frac{\rho_{b}}{\tilde{U}}}\left(\tilde{U}\rho_{b}-{\bm{v}}_{b}^{2}\right),$ (39)
$\mathbf{g}_{ti}=-\sqrt{\frac{\rho_{b}}{\tilde{U}}}{\bm{v}}_{bi},\qquad\mathbf{g}_{ij}=\sqrt{\frac{\rho_{b}}{\tilde{U}}}\delta_{ij},$ (40)
@@
$\sqrt{-\mathbf{g}}=\sqrt{\frac{\rho_{b}^{3}}{\tilde{U}}}.$ (41)
@@
$\mathcal{A}^{\mu}=\Big{(}\frac{f^{\prime\prime}_{3}}{\sqrt{\mathbf{-g}}\tilde{U}},\frac{f^{\prime\prime}_{3}{\bm{v}}_{b}}{\sqrt{\mathbf{-g}}\tilde{U}}\Big{)},$ (42)
@@
$\mathcal{B}^{\mu}=\Big{(}\frac{\ddot{f}_{1}}{\sqrt{\mathbf{-g}}\tilde{U}},\frac{\ddot{f}_{1}{\bm{v}}_{b}}{\sqrt{\mathbf{-g}}\tilde{U}}\Big{)}.$ (43)
@@
$\mathcal{M}^{2}=\frac{1}{\sqrt{\mathbf{-g}}}\Big{[}-\frac{2\ddot{f}_{1}f^{\prime\prime}_{3}}{\tilde{U}}+f^{\prime\prime}_{1}\Big{]}.$ (44)
Note that spacetime indices are raised and lowered throughout with the full metric $\mathbf{g}_{\mu\nu}$, not with the Unruh metric $g_{\mu\nu}$.
@@
The contributions of the final two (vector field) terms in the dynamics of the scalar field are subdominant (are of lower order in $N^{-1/2}$) when compared to the first two terms. The second-order emergent metric $\mathbf{g}_{\mu\nu}$ includes a term $h_{\mu\nu}$ which signifies the deviation from the Unruh paradigm metric $g_{\mu\nu}$. The $N$-scalings of the various terms are as follows
$\begin{split}\mathbf{g}_{\mu\nu}=g_{\mu\nu}[\mathcal{O}(N)]+h_{\mu\nu}[\mathcal{O}(N^{0})],\\ A_{\mu}=[\mathcal{O}(N^{-1})],\\ B_{\mu}=[\mathcal{O}(N^{-2})].\end{split}$ (45)
where $g_{\mu\nu}[\mathcal{O}(N)]$ is the conventional Unruh paradigm metric. Hence to leading order in the large $N$ expansion, corrections to the Unruh metric dominate the quantum nonlinear effects in the Lorentz-invariant wave equation (37).
@@ III.5 Dispersion relation
To understand the nature of the dispersion relation, let us consider a plane wave propagating in $x$ direction (37)
@@
$\omega=v_{bx}k+\frac{1}{2}\left[i(\ddot{f}_{1}-f^{\prime\prime}_{3})\right.\\ \left.\pm\sqrt{4(\tilde{U}f^{\prime\prime}_{1}-\ddot{f}_{1}f^{\prime\prime}_{3}+c^{2}_{s,b}k^{2})-(\ddot{f}_{1}-f^{\prime\prime}_{3})^{2}}\right].$ (47)
Here, we have defined the effective sound velocity of the background ($\rho_{b}$, $\theta_{b}$) to be $c_{s,b}^{2}=\tilde{U}\rho_{b}$. As expected, when $f_{1}=f_{3}=0$, this dispersion relation reduces to the standard Bogoliubov dispersion. The appearance of a generally complex frequency is a direct manifestation of backreaction, causing the condensate particle number to be not separately conserved. The exact rate at which this occurs, and the corresponding correction on the leading-order Bogoliubov dispersion depends on the relevant field configurations leading to the solutions of the BdG equation (6).
@@ IV Quantum nonlinear spacetime metric
Even though the form of the analogue metric that couples to the second-order fluctuation field, $\mathbf{g}_{\mu\nu}$ is formally identical to that of a metric corresponding to the linear regime, the exact functional dependence on the underlying fluid variables is different, due to the fact that now the definition of background has changed. In this section we will explicitly calculate the components of the metric and will see how they differ from the standard Unruh paradigm. In the remainder of the paper we will denote the emergent nonlinear background as background-II (variables $\rho_{b}$, $\theta_{b}$), while the background related to the strict linearization leading to the Unruh paradigm will be denoted without italics and boldface as background-I (variables $\rho_{0}$, $\theta_{0}$). For both backgrounds we demand, as usual, that only perturbations which vary much faster than the background are considered. In particular, the relevant length scales of background-II must always be much less than those of background-I for them to be meaningfully separable.
@@ IV.1 Metric coupled to second-order fluctuations
$\mathbf{g}_{tt}=g_{tt}-\sqrt{\frac{\rho_{0}}{U}}\Big{[}\epsilon U\delta\rho_{1}-2\epsilon{\bm{v}}_{0}\cdot{\bm{v}}_{1}-\epsilon^{2}{\bm{v}_{1}}^{2}\\ +\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}\big{(}c_{s0}^{2}-{\bm{v}}^{2}_{0}+\epsilon U\delta\rho_{1}-2\epsilon{\bm{v}}_{0}\cdot{\bm{v}}_{1}\big{)}-\epsilon^{2}\frac{\delta\rho^{2}_{1}}{8\rho^{2}_{0}}\big{(}c_{s0}^{2}-{\bm{v}}^{2}_{0}\big{)}\Big{]},$ (48)
$\mathbf{g}_{ti}=g_{ti}-\sqrt{\frac{\rho_{0}}{U}}\Big{[}\epsilon{\bm{v}}_{1i}+\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}\big{(}{\bm{v}}_{0}+\epsilon{\bm{v}}_{1}\big{)}_{i}-\epsilon^{2}\frac{\delta\rho^{2}_{1}}{8\rho^{2}_{0}}{\bm{v}}_{0i}\Big{]}.$ (49)
$\mathbf{g}_{ij}=g_{ij}+\sqrt{\frac{\rho_{0}}{U}}\Big{[}\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}-\epsilon^{2}\frac{\delta\rho_{1}^{2}}{8\rho_{0}^{2}}\Big{]}\delta_{ij}.$ (50)
@@
$g_{tt}=-\sqrt{\frac{\rho_{0}}{U}}\left(c_{s0}^{2}-{\bm{v}}_{0}^{2}\right),\qquad g_{ti}=-\sqrt{\frac{\rho_{0}}{U}}{\bm{v}}_{0i},\\ g_{ij}=\sqrt{\frac{\rho_{0}}{U}}\delta_{ij}.$ (51)

### v3

@@ III.1 Linearized fluctuations around a background
and so on. For later convenience, we can thus establish a one-to-one correspondence with the subscript of each component of the density perturbation with the relevant scaling with $N^{1/2}$. For example, the subscript $0\rightarrow\mathcal{O}(N^{1/2})$ and similarly for other terms in descending order, i.e., an index $\alpha$ implies being of ${\cal O}(N^{(1-\alpha)/2})$. Note also that $f_{1}=\mathcal{O}(N^{0})$ and $f_{3}=\mathcal{O}(N^{-1})$ defined earlier in (23) follow the same scaling convention, but that the phase perturbation index in (25) does not indicate a scaling with powers of $N^{1/2}$.
@@
The components of the metric $g_{\mu\nu}$ depend on the background variables $\rho_{0},\theta_{0}$, where the speed of sound in this background is $c^{2}_{s0}=\lambda\rho_{0}$.
@@ III.2 Second-order fluctuations around a background
$\partial_{t}(\rho_{0}+\epsilon\delta\rho_{1})+\nabla\cdot\Big{(}(\rho_{0}+\epsilon\delta\rho_{1})(\nabla\theta_{0}+\epsilon\nabla\delta\theta_{1})\Big{)}=f_{1}(\rho_{0},\theta_{0})+\epsilon\dot{f}_{1}\delta\rho_{1}+\epsilon f_{1}^{\prime}\delta\theta_{1}+\epsilon^{2}\dot{f}_{1}^{\prime}\delta\rho_{1}\delta\theta_{1},$ (28)
$\partial_{t}(\theta_{0}+\epsilon\delta\theta_{1})+\lambda(\rho_{0}+\epsilon\delta\rho_{1})+\frac{1}{2}(\nabla(\theta_{0}+\epsilon\delta\theta_{1}))^{2}+V+f_{3}(\rho_{0},\theta_{0})+\epsilon\dot{f}_{3}\delta\rho_{1}+\epsilon f_{3}^{\prime}\delta\theta_{1}+\epsilon^{2}\dot{f}^{\prime}_{3}\delta\rho_{1}\delta\theta_{1}=0.$ (29)
where we introduced the abbreviating notation that partial derivatives with respect to $\rho_{c}$ and $\theta_{c}$, taken at $\rho_{c}=\rho_{0}$ and $\theta_{c}=\theta_{0}$, are represented as an overdot for $\rho_{c}$ derivatives and as a prime for $\theta_{c}$ derivatives. Note in this regard that an overdot thus implies a scaling factor $1/N$ to leading order. As we will argue later in detail, the two above equations define the new background on which we will consider the propagation of nonlinear fluctuations. Then the new equations governing the perturbations are
$\epsilon^{2}\partial_{t}\delta\rho_{2}+\nabla\cdot\Big{(}\epsilon^{2}\delta\rho_{2}(\nabla\theta_{0}+\epsilon\nabla\delta\theta_{1})+\epsilon^{2}(\rho_{0}+\epsilon\delta\rho_{1})\nabla\delta\theta_{2}\Big{)}=\epsilon^{2}\ddot{f}_{1}\delta\rho_{2}+\epsilon^{2}f_{1}^{\prime\prime}\delta\theta_{2},$ (30)
$\epsilon^{2}\partial_{t}\delta\theta_{2}+\epsilon^{2}\lambda\delta\rho_{2}+\epsilon^{2}\nabla(\theta_{0}+\epsilon\delta\theta_{1})\cdot\nabla\delta\theta_{2}+\epsilon^{2}\ddot{f}_{3}\delta\rho_{2}+\epsilon^{2}f_{3}^{\prime\prime}\delta\theta_{2}=0.$ (31)
@@
where we truncated in the second line the series in Eqs. (24) and (25) after second order in the fluid variables $\delta\rho_{2}$ and $\delta\theta_{2}$. We assume that these combinations satisfy the dynamical fluid Eqs. (21) and (22)) for two different boundary conditions, such that the expansions above are convergent, which in turn requires $\epsilon^{2}\delta\rho_{2}\ll\epsilon\delta\rho_{1}\ll\rho_{0}$. This also implies that we are not separately demanding $\rho_{0},\theta_{0}$ to follow any dynamical equations, at least formally. Then the condition on $\rho_{0},\theta_{0}$ is simply that they are two sufficiently slowly varying “background” functions, which are not necessarily solutions to the fluid dynamical equations, but can be chosen, for example, according to the ease of interpretation of experiments. Namely, such that the first and second order perturbative corrections on top of the background satisfy the fluid-dynamical equations.
@@
Taking into account the variables ($\rho_{b},\theta_{b}$) of Eq. (32) defining the new background, we can substitute $\delta\rho_{2}$ from (31) into (30) to obtain
$\partial_{t}\Big{(}\frac{1}{\tilde{\lambda}}\Big{[}\partial_{t}\delta\theta_{2}+{\bm{v}}_{b}\cdot\nabla\delta\theta_{2}+f^{\prime\prime}_{3}\delta\theta_{2}\Big{]}\Big{)}+\nabla\cdot\Bigg{(}\frac{1}{\tilde{\lambda}}{\bm{v}}_{b}\Big{[}\partial_{t}\delta\theta_{2}+{\bm{v}}_{b}\cdot\nabla\delta\theta_{2}+f^{\prime\prime}_{3}\delta\theta_{2}\Big{]}-\rho_{b}\nabla\delta\theta_{2}\Bigg{)}\\ -\frac{\ddot{f}_{1}}{\tilde{\lambda}}\Big{(}\partial_{t}\delta\theta_{2}+{\bm{v}}_{b}\cdot\nabla\delta\theta_{2}+f^{\prime\prime}_{3}\delta\theta_{2}\Big{)}+f^{\prime\prime}_{1}\delta\theta_{2}=0.$ (34)
We have defined the background velocity
@@
$\tilde{\lambda}=\lambda+\ddot{f}_{3}.$ (36)
@@ III.3 Lorentz-invariant wave equation for second-order perturbations
$\frac{1}{\sqrt{-\mbox{\sf g}}}\mathcal{D}_{\mu}\left(\sqrt{-\mbox{\sf g}}\mbox{\sf g}^{\mu\nu}\mathcal{D}_{\nu}\delta\theta_{2}\right)\\ -\mbox{\sf g}^{\mu\nu}\mathcal{A}_{\mu}\mathcal{D}_{\nu}\delta\theta_{2}-\mathcal{B}^{\mu}\mathcal{D}_{\mu}\delta\theta_{2}+\mathcal{M}^{2}\delta\theta_{2}=0.$ (37)
@@
$\mbox{\sf g}_{tt}$ $=$ $-\sqrt{\frac{\rho_{b}}{\tilde{\lambda}}}\left(\tilde{\lambda}\rho_{b}-{\bm{v}}_{b}^{2}\right),$ (39)
$\mbox{\sf g}_{ti}$ $=$ $-\sqrt{\frac{\rho_{b}}{\tilde{\lambda}}}{\bm{v}}_{bi},\qquad\mbox{\sf g}_{ij}=\sqrt{\frac{\rho_{b}}{\tilde{\lambda}}}\delta_{ij},$ (40)
@@
$\sqrt{-\mbox{\sf g}}=\sqrt{\frac{\rho_{b}^{3}}{\tilde{\lambda}}}.$ (41)
@@
$\mathcal{A}^{\mu}=\Big{(}\frac{f^{\prime\prime}_{3}}{\sqrt{\mbox{$-$\sf g}}\tilde{\lambda}},\frac{f^{\prime\prime}_{3}{\bm{v}}_{b}}{\sqrt{\mbox{$-$\sf g}}\tilde{\lambda}}\Big{)},$ (42)
@@
$\mathcal{B}^{\mu}=\Big{(}\frac{\ddot{f}_{1}}{\sqrt{\mbox{$-$\sf g}}\tilde{\lambda}},\frac{\ddot{f}_{1}{\bm{v}}_{b}}{\sqrt{\mbox{$-$\sf g}}\tilde{\lambda}}\Big{)}.$ (43)
@@
$\mathcal{M}^{2}=\frac{1}{\sqrt{\mbox{$-$\sf g}}}\Big{[}-\frac{2\ddot{f}_{1}f^{\prime\prime}_{3}}{\tilde{\lambda}}+f^{\prime\prime}_{1}\Big{]}.$ (44)
Note that spacetime indices are raised and lowered throughout with the full metric $\mbox{\sf g}_{\mu\nu}$, not with the Unruh metric $g_{\mu\nu}$.
@@
The contributions of the final two (vector field) terms in the dynamics of the scalar field are subdominant (are of lower order in $N^{-1/2}$) when compared to the first two terms. The second-order emergent metric $\mbox{\sf g}_{\mu\nu}$ includes a term $h_{\mu\nu}$ which signifies the deviation from the Unruh paradigm metric $g_{\mu\nu}$. The $N$-scalings of the various terms are, to leading order, as follows
$\begin{split}\mbox{\sf g}_{\mu\nu}=g_{\mu\nu}[\mathcal{O}(N)]+h_{\mu\nu}[\mathcal{O}(N^{0})],\\ A_{\mu}=[\mathcal{O}(N^{-1})],\\ B_{\mu}=[\mathcal{O}(N^{-2})],\\ \mathcal{M}^{2}=[\mathcal{O}(N^{-1})],\end{split}$ (45)
where $g_{\mu\nu}[\mathcal{O}(N)]$ is the conventional Unruh paradigm analogue metric. Hence to leading order in the large $N$ expansion, corrections to the Unruh metric dominate the quantum nonlinear effects in the Lorentz-invariant wave equation (37).
@@ III.5 Dispersion relation
To understand the nature of the dispersion relation, let us consider a plane wave solution of (37) propagating in $x$ direction,
@@
$\omega=v_{bx}k+\frac{1}{2}\left[i(\ddot{f}_{1}-f^{\prime\prime}_{3})\right.\\ \left.\pm\sqrt{4(\tilde{\lambda}f^{\prime\prime}_{1}-\ddot{f}_{1}f^{\prime\prime}_{3}+c^{2}_{s,b}k^{2})-(\ddot{f}_{1}-f^{\prime\prime}_{3})^{2}}\right].$ (47)
Here, we have defined the effective sound velocity of the background ($\rho_{b}$, $\theta_{b}$) to be $c_{s,b}^{2}=\tilde{\lambda}\rho_{b}$. As expected, when $f_{1}=f_{3}=0$, this dispersion relation reduces to the standard Bogoliubov dispersion. The appearance of a generally complex frequency is a direct manifestation of backreaction, causing the condensate particle number to be not separately conserved. The exact rate at which this occurs, and the corresponding correction on the leading-order Bogoliubov dispersion depends on the relevant field configurations leading to the solutions of the BdG equation (6).
@@ IV Quantum nonlinear spacetime metric
Even though the form of the analogue metric that couples to the second-order fluctuation field, $\mbox{\sf g}_{\mu\nu}$ is formally identical to that of a metric corresponding to the linear regime, the exact functional dependence on the underlying fluid variables is different, due to the fact that now the definition of background has changed. In this section we will explicitly calculate the components of the metric and will see how they differ from the standard Unruh paradigm. In the remainder of the paper we will denote the emergent nonlinear background as background-II (variables $\rho_{b}$, $\theta_{b}$), while the background related to the strict linearization leading to the Unruh paradigm will be denoted without italics and boldface as background-I (variables $\rho_{0}$, $\theta_{0}$). For both backgrounds we demand, as usual, that only perturbations which vary much faster than the background are considered. In particular, the relevant length scales of background-II must always be much less than those of background-I for them to be meaningfully separable.
@@ IV.1 Metric coupled to second-order fluctuations
$\mbox{\sf g}_{tt}=g_{tt}-\sqrt{\frac{\rho_{0}}{\lambda}}\Big{[}\epsilon\lambda\delta\rho_{1}-2\epsilon{\bm{v}}_{0}\cdot{\bm{v}}_{1}-\epsilon^{2}{\bm{v}_{1}}^{2}\\ +\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}\big{(}c_{s0}^{2}-{\bm{v}}^{2}_{0}+\epsilon\lambda\delta\rho_{1}-2\epsilon{\bm{v}}_{0}\cdot{\bm{v}}_{1}\big{)}-\epsilon^{2}\frac{\delta\rho^{2}_{1}}{8\rho^{2}_{0}}\big{(}c_{s0}^{2}-{\bm{v}}^{2}_{0}\big{)}\Big{]},$ (48)
$\mbox{\sf g}_{ti}=g_{ti}-\sqrt{\frac{\rho_{0}}{\lambda}}\Big{[}\epsilon{\bm{v}}_{1i}+\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}\big{(}{\bm{v}}_{0}+\epsilon{\bm{v}}_{1}\big{)}_{i}-\epsilon^{2}\frac{\delta\rho^{2}_{1}}{8\rho^{2}_{0}}{\bm{v}}_{0i}\Big{]}.$ (49)
$\mbox{\sf g}_{ij}=g_{ij}+\sqrt{\frac{\rho_{0}}{\lambda}}\Big{[}\epsilon\frac{\delta\rho_{1}}{2\rho_{0}}-\epsilon^{2}\frac{\delta\rho_{1}^{2}}{8\rho_{0}^{2}}\Big{]}\delta_{ij}.$ (50)
@@
$g_{tt}=-\sqrt{\frac{\rho_{0}}{\lambda}}\left(c_{s0}^{2}-{\bm{v}}_{0}^{2}\right),\qquad g_{ti}=-\sqrt{\frac{\rho_{0}}{\lambda}}{\bm{v}}_{0i},\\ g_{ij}=\sqrt{\frac{\rho_{0}}{\lambda}}\delta_{ij}.$ (51)